OpenStudy (anonymous):

P = L [(r/12) / (1 - (1 + r/12) ^ (-t)] P = monthly payment L = loan amount r = annual rate of interest, expressed as a decimal t= length of loan, in months On July 8, 2003, average interest rates were: (a) 5.52% on 30-year mortgages (b) 4.85% on 15-year mortgages 1. For each rate, calculate the monthly payment for a loan of $200,000. 2. Then compute the total amount paid over the term of each loan. 3. Calculate the interest paid on each loan.

OpenStudy (anonymous):

Would the numbers towards the end (1-3) be the steps to get my answer or are they separate problems for the given (a) and (b)?

OpenStudy (ahsome):

\[P=L((r/12)/(1-(1+r/12)^-t)\] Rewrote equation for others to see easier

OpenStudy (anonymous):

\[P =L \left[ \frac{ \frac{ r }{ 12 } }{ 1-(1+\frac{ r }{ 12 })^{-t} } \right]\]

OpenStudy (anonymous):

Is that better? I also have the labels for the variables in the question I opened :) Also, would the numbers towards the end (1-3) be the steps to get my answer or are they separate problems for the given (a) and (b)?

OpenStudy (ahsome):

Because of (a) and (b), you would have to do for both rates, unless it says use the 15 year or 30 year rate

OpenStudy (aum):

1 to 3 are three separate questions. For (1), plug the values into the formula and calculate the monthly payment P. For (2), Multiply P calculated above by 30 * 12 (for 30 year mortgage) or P * 15 * 12 (for 15 year mortgage) to find the total amount paid. For (3), Subtract the loan amount of $200,000 from the total amount paid calculated in step 2 to find the total interest paid.

OpenStudy (anonymous):

@aum It makes much more sense now! Thank you very much! :D

OpenStudy (aum):

You are welcome.